A MOTHER IS 6 TIMES AS OLD AS HIS SON AND THREE TIMES AS OLD AS HIS DAUGHTER. IF THE DAUGHTER IS 5 YEARS OLDER THAN THE



Answer by ankor@dixie-net.com(22740)

A MOTHER IS 6 TIMES AS OLD AS HIS SON AND THREE TIMES AS OLD AS HIS DAUGHTER.
IF THE DAUGHTER IS 5 YEARS OLDER THAN THE son. Then how old are they all now
:
Mother is 6 times as old as son
m = 6s
Mother is 3 times as old as daughter
m = 3d
Daughter is 5 yrs old than son
d = (s+5)
:
m = 3d
replace d with (s+5)
m = 3(s+5)
m = 3s + 15
:
m = 6s
replace m with (3s+15)
3s + 15 = 6s
15 = 6s - 3s
15 = 3s
s = 15/3
s = 5 yrs old is the son
:
I'll let you find the ages of Mom and daughter.
Check your solutions in the given statements